### v1

# DynaSpec: Context-aware Dynamic Speculative Sampling for Large-Vocabulary Language Models

Abstract
Speculative decoding (a.k.a. speculative sampling) has become a standard way to accelerate LLM inference: a small drafter proposes multiple tokens and a large target model verifies them once per speculation length. Recently, scaling of the LLM vocabulary has pushed the number of tokens to grow substantially. While verification over the full vocabulary leaves the target model largely unaffected, the $\mathcal{O}(|V|d)$ parameters in the drafter’s output head become a latency bottleneck, slowing the entire pipeline. Contemporary methods (e.g., FR-Spec, VocabTrim) restrict the drafter’s vocabulary to a fixed subset of the target model’s vocabulary, ranked in descending order of token frequency. Although this reduces draft-time compute, it is brittle, since : (i) frequency lists are corpus-dependent and require retuning to generalize, and (ii) static shortlists suppress rare or domain-specific tokens, lowering the expected number of tokens per verification step. We propose DynaSpec, a context-dependent dynamic shortlisting mechanism that is robust, speeds up drafting, and generalizes across diverse tasks. Concretely, we introduce lightweight, coarse-grained meta-classifiers that route contexts to a small number of token clusters; the union of the top-$k$ selected clusters forms the drafter’s shortlist, while verification retains the full vocabulary and exactness. The meta-classifier finishes its computation earlier than the drafter’s hidden state generation by exploiting parallel execution of draft encoding and meta shortlisting on separate streams. On standard speculative-decoding benchmarks, we observe consistent gains in mean accepted length over fixed-shortlist baselines, while context-dependent selection enables smaller shortlists without degrading acceptance.
1 Introduction
Over the past few years, large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023; Liu et al., 2024; Jiang et al., 2024) have improved dramatically, reshaping products and workflows across sectors. Beyond algorithmic advances, much of this progress has come from scaling data and especially model parameters (Kaplan et al., 2020; Hoffmann et al., 2022). Yet these gains come with a cost: inference becomes compute-heavy and latency-sensitive, which limits real-time applications (Kwon et al., 2023; Yuan et al., 2024). As agentic systems (Yao et al., 2023; Wang et al., 2024) and reasoning-heavy workflows (e.g., chain-of-thought) (Wei et al., 2022; Zhang et al., 2024) proliferate, the demand for faster inference keeps rising. Speculative decoding (Chen et al., 2023; Leviathan et al., 2023; Li et al., 2024a, b) is a practical way to accelerate generation. A smaller draft model proposes multiple tokens so the target model need not run at every step, while the rejection sampling procedure guarantees that the final output distribution matches that of the target model. In the worst case, quality and latency revert to standard decoding; in typical cases, throughput improves substantially.
Recently scaling laws for vocabulary (Tao et al., 2024; Huang et al., ; Yu et al., 2025) suggests that larger models benefit from larger vocabularies, and deployed LLMs have indeed expanded their vocabularies over time. For example, tokenizer vocabularies have scaled from 32k entries in Llama-2 (Touvron et al., 2023) to 128k in Llama-3 (Grattafiori et al., 2024), 129k in DeepSeek-V3 (Liu et al., 2024), 152k in Qwen-2.5 (Yang et al., 2025), and 262k in Gemma-3 (Team et al., 2025). While the large vocabulary size contributes only marginally to the inference cost of the target model’s output layers given its overall model size, it imposes a substantial computational burden on the draft model during speculative decoding, as the draft model is much smaller and the output layer constitutes a larger fraction of its total computation. thereby diminishing the overall pipeline efficiency. To mitigate this, FR-Spec (Zhao et al., 2025) and Vocabtrim (Goel et al., 2025) replace the full vocabulary with a fixed shortlist—typically the top-p% of the most frequent tokens measured on a reference corpus. This speeds up drafting, but it is suboptimal for two reasons: (1) frequency lists are corpus-dependent and often fail to generalize across benchmarks, requiring substantial retuning; and (2) a static subset of the token vocabulary can degrade performance on tasks that rely on rare or domain-specific tokens, where maintaining diversity in the candidate set is critical for high-quality outputs.
In this work, we move beyond fixed shortlists and introduce a dynamic vocabulary head. We first define meta-labels as coarse token clusters: the vocabulary is partitioned into $M\!\ll\!|V|$ clusters $\{C_{m}\}_{m=1}^{M}$ by clustering column-normalized LM-head weights. Conditioned on embedding and features, a lightweight meta-classifier scores clusters and we evaluate the drafter’s LM head only on the union of the selected clusters $\mathcal{V}_{S}=\bigcup_{m\in\mathcal{K}}\!C_{m}$, while verification still runs over the full vocabulary, preserving exactness. This coarse-to-fine routing, inspired by Jiang et al. (2021) and  Kharbanda et al. (2022), replaces an $\mathcal{O}(|V|d)$ head with $\mathcal{O}\!\big((M+|\mathcal{V}_{S}|)d\big)$. Making the support context-dependent leads to a consistent increase (across datasets) in the mean accepted length relative to the static subset proposed in Zhao et al. (2025). However, a dynamic (indexed) head introduces gathered matmul overhead and can be slower than a fixed, pre-trimmed dense head. We therefore couple the dynamic head with a position-aware cluster budget: early tokens receive larger shortlists, and the budget decays with position. As argued in recent works, Li et al. (2025, 2024b), prioritizing earlier tokens improves expected acceptance, while shrinking the shortlist later reduces latency, yielding net end-to-end gains without altering verification. Finally, we show the same position-aware schedule can retrofit frequency-ranked shortlists (FR-Spec) (Zhao et al., 2025); see Appendix A.1 for details.
Contributions.
To reconcile acceptance gains from a dynamic head with the matmul overhead it introduces, DynaSpec combines context-aware routing with a position-aware budget. Concretely: (i) we introduce meta-labels (coarse token clusters) and a lightweight meta-classifier that selects a context-dependent candidate shortlist of clusters on which the drafter operates, while verification remains exact on the full vocabulary; (ii) we show that dynamic, context-conditioned supports dominate any static subset in expected acceptance; (iii) we mitigate indexed-head latency via a position-aware schedule $k_{\mathrm{c}}(t)$, allocating more candidate clusters early and fewer later, with a fused index+GEMM kernel, yielding lower draft-time compute without sacrificing early-step acceptance; and (iv) empirically, across seven tasks and LLaMA models with $128\mathrm{k}$ vocabularies, DynaSpec increases mean accepted length over fixed-shortlist baselines. The approach is plug-in compatible with EAGLE-style pipelines and orthogonal to system-side optimizations, making it a practical drop-in for large-vocabulary deployments.
2 Related Work
2.1 LLMs with Large Vocabulary
As LLMs have grown in scale, so too have their vocabularies. Early models such as GPT-3 (Brown et al., 2020) and LLaMA-2 (Touvron et al., 2023) used vocabularies of $32\mathrm{k}$ tokens, while more recent releases exceed $100\mathrm{k}$ (e.g., $128\mathrm{k}$ in LLaMA-3 (Grattafiori et al., 2024), $152\mathrm{k}$ in Qwen-2.5 (Yang et al., 2025), and $262\mathrm{k}$ in Gemma-3 (Team et al., 2025)). Large vocabularies are especially common in multilingual and high-coverage LLMs: mT5 (Xue et al., 2020) adopts a $250\mathrm{k}$ SentencePiece (Kudo and Richardson, 2018) vocabulary to improve cross-lingual coverage, and PaLM uses a $256\mathrm{k}$ unigram vocabulary to shorten sequences and better encode code and symbols (Chowdhery et al., 2023). Recently, empirical scaling laws suggest that larger vocabularies improve expressivity and perplexity (Tao et al., 2024; Yu et al., 2025), but they also magnify inference cost: the output projection grows linearly with vocabulary size, which can dominate latency in small drafters.
2.2 Speculative Decoding
In line with Speculative Decoding, early work accelerates greedy generation (Stern et al., 2018; Sun et al., 2021), and subsequent work extends the idea to non-greedy sampling with distributional guarantees (Leviathan et al., 2023; Chen et al., 2023; Xia et al., 2022). These methods differ primarily in how they form proposals: some reuse information already present in the prompt via retrieval-guided drafting—effective when outputs strongly overlap inputs (Saxena, 2023; Yang et al., 2023; He et al., 2024)—whereas others learn parametric drafters for general-purpose speedups, ranging from parallel MLP heads (and serial variants) to lightweight transformer blocks that condition on richer context (Cai et al., 2024; Ankner et al., 2024). The EAGLE series further refines this paradigm: EAGLE employs lightweight transformer drafters, while EAGLE-2 introduces dynamic draft trees to boost efficiency (Li et al., 2024a, b). Orthogonal advances in training and systems—such as adaptive exploration, cache reuse, and efficient serving stacks—increase accepted tokens per verification and reduce overhead (Kwon et al., 2023; Zheng et al., 2024; Dao, 2023). Complementing these directions, MagicDec analyzes the latency–throughput trade-off and shows that speculative decoding remains effective at large batch sizes and with long contexts, whereas TriForce targets long-context regimes by compressing the drafter’s KV cache and coordinating reuse across draft steps (Sadhukhan et al., 2024; Sun et al., 2024). Most recently, Timor et al. (2025) proposes lossless speculative decoding for heterogeneous vocabularies, relaxing the same-tokenizer requirement between drafter and target; meanwhile, Li et al. (2025) combines Medusa-style heads (Cai et al., 2024) with the EAGLE (Li et al., 2024b) framework. We target a different bottleneck: as vocabularies expand, the drafter’s output projection increasingly dominates latency.
2.3 Accelerating Large Vocabularies
The computational bottleneck of large output spaces or vocabularies has been recognized in both language modeling and neighboring fields. Within LLMs, FR-Spec (Zhao et al., 2025) and VocabTrim (Goel et al., 2025) mitigate drafter overhead by pruning the output layer to a fixed shortlist of high-frequency tokens. While effective in reducing latency, these static subsets may suppress rare or domain-specific tokens, lowering acceptance in speculative decoding. Other work on large vocabulary LLM training, such as CCE (Wijmans et al., 2024), reduces peak memory by fusing cross-entropy directly into the kernel, avoiding explicit logit materialization. Our approach is motivated by research in extreme classification, where the goal is to accelerate learning and inference over output spaces with millions of labels. Methods such as LightXML (Jiang et al., 2021), CascadeXML (Kharbanda et al., 2022), and related meta classifier based strategies (Zhang et al., 2021; Kharbanda et al., 2023) adopt coarse-to-fine or cluster-based mechanisms that limit computation to a small subset of relevant labels per input. These ideas directly inspire DynaSpec: rather than evaluating the drafter across the entire vocabulary or a static shortlist, we employ a lightweight meta-classifier to select a small number of context-relevant clusters, retaining the coverage of rare tokens while reducing compute.
3 Theoretical Analysis
Let $T_{T}$ denote the per-token latency of the target model under standard autoregressive decoding. In speculative decoding, let $T_{D}$ be the draft model’s per-token generation time, and let $T_{V}(\gamma)$ be the time for the target to verify $\gamma$ proposed tokens. Given $\alpha\in[0,1]$ as the expected acceptance rate and $\gamma\in\mathbb{N}$ the speculation length, following standard derivations (Leviathan et al., 2023), the expected number of tokens produced per verification step is
$\Omega(\alpha,\gamma)=\mathbb{E}[\#\text{generated tokens}]=\frac{1-\alpha^{\gamma+1}}{1-\alpha}.$ (1)
The average per-token latency under speculative decoding is therefore
$T_{\mathrm{SD}}=\frac{\gamma\cdot T_{D}+T_{V}(\gamma)}{\Omega(\alpha,\gamma)}.$
Hence, the speedup relative to standard autoregressive decoding is
$\mathrm{Speedup}=\frac{T_{T}}{T_{\mathrm{SD}}}=\frac{T_{T}\cdot\Omega(\alpha,\gamma)}{\gamma\cdot T_{D}+T_{V}(\gamma)}.$ (2)
3.1 Large-Vocabulary Bottleneck on Speedup
As shown in Equation 2 the speedup of speculative decoding is determined by the verification time $T_{V}(\gamma)$, the draft cost $T_{D}$, and the expected generation length $\Omega(\alpha,\gamma)$. As the tokenizer vocabulary grows, the target-side terms $T_{T}$ and $T_{V}(\gamma)$ are comparatively insensitive to $|V|$, since their per-token runtime is dominated by stacked attention and MLP compute across many layers. In contrast, the drafting time per-token decomposes into a transformer core term $\Theta(L_{D}d^{2}+L_{D}Sd)$ (attention and MLP) and a vocabulary-head term $\Theta(|V|d)$. As $V$ increases, the $\Theta(|V|d)$ component can overtake the core, inflating $T_{D}$ and—because it is multiplied by $\gamma$ in the speedup denominator—either reducing the achievable speedup or forcing a smaller $\gamma$.
Restricting the draft to a smaller subset of the vocabulary effectively reduces the $\Theta(|V|d)$ term and can lower $T_{D}$. However, if this subset is not context-aware, the acceptance rate $\alpha$ may drop, shrinking $\Omega(\alpha,\gamma)$ and hence offsetting the latency gains, as we discuss next.
3.2 Impact of Static Sub-Vocabulary Constraints on Acceptance
 [^1] For details, follow section 3 in (Leviathan et al., 2023)
We use context $c$ interchangeably with the conditioning prefix $x_{<t}$. Assuming $p_{c}(x)\;\equiv\;p(x_{t}{=}x\mid x_{<t}{=}c)$, the target distribution and $q_{c}(x)\;\equiv\;q(x_{t}{=}x\mid x_{<t}{=}c),$ for the draft model over the full vocabulary $V$. The acceptance probability (acceptance rate) is
$\beta(c;q_{c})\;=\;\sum_{x\in V}\min(p_{c}(x),\,q_{c}(x))\;=\;\mathbb{E}_{x\sim q_{c}}\!\left[\min\!\Big(1,\tfrac{p_{c}(x)}{q_{c}(x)}\Big)\right].$ (3)
Under the standard i.i.d. assumption for $\beta$ , the average acceptance is
$\alpha\;:=\;\mathbb{E}_{c}\!\left[\beta(c;q_{c})\right].$ (4)
Lemma 1 (Acceptance on a fixed support). For any context $c$ and any subset $S\subseteq V$ with $|S|=k$, following Equation 3, the acceptance rate can be written as
$A^{\!*}(c;S)\;:=\;\max_{q_{c}:\ \mathrm{supp}(q_{c})\subseteq S}\ \sum_{x\in V}\min\{p_{c}(x),q_{c}(x)\}\;=\;\sum_{x\in S}p_{c}(x)\,$
That is, $A^{\!*}(c;S)$ is the maximum of $\beta(c;q_{c})$ over all $q_{c}$ whose support is contained in $S$.
Proof. For $x\notin S$, the overlap $\min\{p_{c}(x),q_{c}(x)\}=0$. For $x\in S$, $\min\{p_{c}(x),q_{c}(x)\}\leq p_{c}(x)$ with equality iff $q_{c}(x)\geq p_{c}(x)$. Normalizing $p_{c}$ on $S$ makes this hold for all $x\in S$. Summing over $x\in S$ gives the value. $\square$
Lemma 2 (Oracle or Pointwise optimal subset). Let $p_{c}^{\downarrow}(1)\geq\cdots\geq p_{c}^{\downarrow}(|V|)$ be the sorted probabilities and define the top-$k$ mass $H_{k}(p_{c}):=\sum_{i=1}^{k}p_{c}^{\downarrow}(i)$. Then
$\max_{|S|=k}A^{\!*}(c;S)\;=\;H_{k}(p_{c}),$
attained by any top-$k$ subset under $p_{c}$ (ties allowed).
Proof. By Lemma 1, $\max_{|S|=k}A^{\!*}(c;S)=\max_{|S|=k}\sum_{x\in S}p_{c}(x)=H_{k}(p_{c})$ (rearrangement/majorization). $\square$
Theorem A (Oracle dynamic dominates any static subset). Let $S_{c}^{\star}$ be the oracle top-$k$ set for context $c$, i.e., any measurable top-$k$ subset under $p_{c}$. For any static baseline $S_{\mathrm{stat}}$ with $|S_{\mathrm{stat}}|=k$,
$\mathbb{E}_{c}\!\big[A^{\!*}(c;S_{c}^{\star})\big]\;=\;\mathbb{E}_{c}\!\big[H_{k}(p_{c})\big]\;\geq\;\mathbb{E}_{c}\!\big[A^{\!*}(c;S_{\mathrm{stat}})\big]\;=\;\mathbb{E}_{c}\!\Big[\sum_{x\in S_{\mathrm{stat}}}p_{c}(x)\Big],$
with strict inequality whenever, on a set of contexts of positive probability, the (unique) top-$k$ set differs from $S_{\mathrm{stat}}$.
Proof. Pointwise, $A^{\!*}(c;S_{c}^{\star})=H_{k}(p_{c})\geq\sum_{x\in S_{\mathrm{stat}}}p_{c}(x)=A^{\!*}(c;S_{\mathrm{stat}})$ by Lemma 2. Take expectations; strictness follows from a strict pointwise gap on a set of positive measure. $\square$
Consequence (expected tokens per verification). Fix speculation length $\gamma$. Since $\Omega(\alpha,\gamma)$ is strictly increasing in $\alpha$, Theorem A implies
$\Omega\!\big(\alpha_{\mathrm{dyn}}^{\star},\gamma\big)\;\geq\;\Omega\!\big(\alpha_{\mathrm{stat}},\gamma\big),$
where $\alpha_{\mathrm{dyn}}^{\star}$ denotes the optimal expected acceptance rate. A dynamic context-aware subset yields at least as many, and typically more generated tokens per verification step by retaining more target probability mass within the drafting support than any static subset.
In the next section, we propose DynaSpec, a context-aware, meta-classifier–driven dynamic vocabulary head that reduces the drafter’s per-token cost $T_{D}$ while maintaining a high $\Omega(\alpha,\gamma)$; verification continues over the full vocabulary, preserving exactness.
4 DynaSpec
Unlike a fixed output head, DynaSpec selects tokens dynamically based on context. It uses a lightweight meta-classifier (the router) to route hidden state of draft model to coarse token clusters, decoding only over the resulting shortlist of tokens within the clusters that are selected by the meta-classifier.
4.1 Preliminary
Target LLM. Let $\mathcal{T}$ represent the target LLM over vocabulary $\mathcal{V}$. Given an input sequence $x_{1:n}$, the embedding layer produces
$Z_{0}=\mathbf{E}(x_{1:n})\in\mathbb{R}^{n\times d}.$
Applying $L$ transformer blocks $\mathcal{F}_{1},\ldots,\mathcal{F}_{L}$ yields
$H\;=\;(\mathcal{F}_{L}\circ\cdots\circ\mathcal{F}_{1})(Z_{0})\in\mathbb{R}^{n\times d}.$
An LM head with weights $W_{LM}\in\mathbb{R}^{d\times|\mathcal{V}|}$ maps to logits $O=HW_{LM}$, and the token probabilities are
$p_{\mathcal{T}}(y_{i}\mid x_{\leq i})\;=\;\operatorname{softmax}(O_{i}).$
Full Vocabulary Drafter. Speculative decoding methods (more specifically EAGLE (Li et al., 2024a, b) series of models) train a lightweight draft model $\mathcal{D}$ (one transformer block, same latent width $d$) to approximate $\mathcal{T}$ at a lower cost. $\mathcal{D}$ reuses and freezes the embedding $\mathbf{E}$ and LM head $W_{LM}$ from $\mathcal{T}$; only its single transformer block is trained:
$\tilde{H}\;=\;\mathcal{D}\!\big(\mathbf{E}(x_{1:n})\big),\qquad\tilde{p}_{\text{EAGLE}}\;=\;\operatorname{softmax}(\tilde{H}W_{LM}).$
Static Sub-vocabulary Drafter. FR-Spec (Zhao et al., 2025) uses a frequency-ranked drafting with vocabulary $\mathcal{V}_{\text{high}}\!\subset\!\mathcal{V}$ the subset of high-frequency tokens (from corpus statistics). Concretely, select the corresponding LM-head columns.
$\widetilde{W}\in\mathbb{R}^{d\times|\mathcal{V}_{\text{high}}|},\quad\widetilde{W}[:,j]=W_{LM}[:,\,\mathcal{V}_{\text{high}}[j]],$
and compute
[…]
4.2 Meta classifiers: Context-aware token Selection
We use a lightweight meta-classifier (interchangeably, the router) to select coarse token clusters (“meta-classes”) per context; details follow.
Vocabulary partition.
We partition the vocabulary $\mathcal{V}$ into $M$ coarse token clusters $\Pi\triangleq\{C_{1},\ldots,C_{M}\}$ via a clustering map $\tau:\mathcal{V}\!\to\![M]$. Following extreme-classification style routing, clusters are formed offline by spherical $k$-means on column-normalized LM-head weights $\{W_{LM}[:,v]/\|W_{LM}[:,v]\|_{2}\}_{v\in\mathcal{V}}$. We do not enforce balance (i.e., $|C_{m}|$ may vary), which empirically preserves recall for semantically dense regions.
Low-cost router.
Given the token embedding $\mathbf{E}(x_{t})$ and the drafter’s previous hidden features $\tilde{H}_{t-1}$, a lightweight MLP $r_{\theta}:\mathbb{R}^{d_{r}}\!\to\!\mathbb{R}^{M}$ produces scores $s=r_{\theta}([\mathbf{E}(x_{t}),\tilde{H}_{t-1}])$ over clusters. The router runs on a separate parallel CUDA stream and completes while the drafter’s attention/MLP is executing, making its wall-clock overhead negligible.
Position-aware cluster selection and shortlist construction.
Instead of a fixed top-$k$ policy, we use a position-aware budget $k_{\mathrm{c}}(t)\in\{k_{\min},\ldots,k_{\max}\}$ that allocates more clusters to early tokens and fewer to later ones. Concretely, with $t=0$ denoting the first drafted token, we keep the first two steps at the full budget and then decay as
$k_{\mathrm{c}}(t)=\begin{cases}k_{\max},&t\in\{0,1\},\\[4.0pt] \left\lfloor\dfrac{k_{\max}}{(t+1)\cdot 2}\right\rfloor,&t\geq 2.\end{cases}$
Given router scores $s$, we select $\mathcal{K}(c,t)=\operatorname{TopK}_{k_{\mathrm{c}}(t)}(s)$ and construct the shortlist $\mathcal{V}_{S}(c,t)=\bigcup_{m\in\mathcal{K}(c,t)}C_{m}$.
Training the router.
We cast routing as multi-label prediction over $M$ clusters with hard labels. Given the hidden state $\tilde{H}_{t}$ of the drafter, let $S^{+}$ denote the positive samples from the drafter model with full vocabulary at draft step $t$ (e.g., the top-$L$ tokens at step $t$). A cluster is positive iff it contains at least one positive token:
$y_{m}\;=\;\mathbb{I}\!\big[C_{m}\cap S^{+}\neq\varnothing\big]\in\{0,1\}.$
We train with binary cross-entropy over all meta labels where $\hat{y}_{m}=\sigma\!\big(s_{m})\big)$:
$\mathcal{L}_{\mathrm{router}}\;=\;\mathbb{E}\!\left[\sum_{m=1}^{M}\Big(-\,y_{m}\log\hat{y}_{m}-(1-y_{m})\log\big(1-\hat{y}_{m}\big)\Big)\right].$
Figure 1: Draft process of DynaSpec. The backbone structure is based on EAGLE where the meta classifier selects the top-k tokens dynamically based on the context.
Algorithm 1 DynaSpec draft steps (single verification cycle)
0: prefix sequence $c$, speculation length $\gamma$, target model $\mathcal{T}$, router $r_{\theta}$, drafter $\mathcal{D}$, LM head $W_{LM}$, clusters $\{C_{m}\}$ over vocabulary, first token cluster budget $k_{\mathrm{max}}$, token budget $k_{\mathrm{t}}$, empty draft list $d$, empty draft list score d_scoresd\_scores
1:for $j=0$ to $\gamma$ do
2:  if $j=0$ then
3:   (stream $S_{m}$) $s\leftarrow r_{\theta}([h_{c[-1]},emb(x_{j})])$; $\mathcal{K}\leftarrow\operatorname{TopK}_{k_{\mathrm{max}}}(s)$; $I\leftarrow\mathrm{indices}\!\big(\cup_{m\in\mathcal{K}}C_{m}\big)$
4:   (stream $S_{d}$) $\tilde{h}\leftarrow\mathcal{D}\!\big([h_{c},c[1:]+x_{j}]\big)$
5:   $last\_step\_scores\leftarrow 0$
6:  else
7:   if $j>1$ then $k_{\mathrm{c}}\leftarrow k_{\mathrm{max}}//((i+1)*2)$ else $k_{\mathrm{c}}\leftarrow k_{\mathrm{max}}$
8:   (stream $S_{m}$) $s\leftarrow r_{\theta}([h_{j-1},emb(x_{j})])$; $\mathcal{K}\leftarrow\operatorname{TopK}_{k_{\mathrm{c}}}(s)$; $I\leftarrow\mathrm{indices}\!\big(\cup_{m\in\mathcal{K}}C_{m}\big)$
9:   (stream $S_{d}$) $\tilde{h}\leftarrow\mathcal{D}\!\big([h_{j-1},x_{j}]\big)$
10:  end if
11:  sync $S_{m},S_{d}$; $z\leftarrow\mathrm{FUSED\_INDEX\_GEMM}\!\big(\tilde{h},\ W_{LM},\ I\big)$ // shortlist logits on gathered columns II
12:  $p\leftarrow\operatorname{log\_softmax}(z)$;  $T_{j},TopP_{j}\leftarrow\operatorname{TopK}_{k_{\mathrm{t}}}(p)$;  $\tilde{T_{j}}=remap2realid(T_{j})$
13:  $cu\_scores\leftarrow TopP_{j}+last\_step\_scores$
14:  $d\leftarrow d\oplus\tilde{T}_{j}$$d\_score\leftarrow d\_score\oplus cu\_scores$ // append score and token to the draft list
15:  $TopC_{j},last\_step\_scores\leftarrow TopK_{k_{\mathrm{t}}}(cu\_scores)$
16:  $x_{j}\leftarrow\tilde{T}_{j}[TopC_{j}]$
17:  if $j=0$ then $h_{j}\leftarrow\tilde{h}[-1]$ else $h_{j}\leftarrow\tilde{h}[TopC_{j}]$
18:end for
19: Re-ranking the $d$ based on their corresponding scores d_scoresd\_scores
20: verify once with $\mathcal{T}$ over full $\mathcal{V}$ on the top-$k$ subset of $d$;
4.3 Dynamic Drafting
At each step $t$, the router (meta-classifier) selects the shortlist indices, while the drafter generates the hidden state for that step. A single gathered-head GEMM is then performed over the shortlisted tokens using this hidden state. The router runs on a separate CUDA stream and typically finishes before the drafter block. The total draft time for a single pass can be expressed as,
$T_{D}(c,t)\ \approx\ T_{\text{embed}}+\max\!\big\{T_{\text{core}},\,T_{\text{meta}}\big\}\ +T_{\text{index+gemm}}\!\big(B(c,t)\big),$
where $T_{\text{meta}}\!\ll\!T_{\text{core}}$ and $T_{\text{embed}}$ is generally small. Thus the only vocabulary-dependent term is the $d\times B(c,t)$ head compute (with $B(c,t)\!\ll\!|\mathcal{V}|$), reducing the denominator in Eq. 2 while preserving acceptance via context-aware support. Since $B(c,t)\!\propto\!|\mathcal{V}_{S}(c,t)|$ and the position-aware budget $k_{\mathrm{c}}(t)$ decay with $t$, the shortlist and thus the head compute shrink as generation proceeds, while larger early budgets preserve acceptance. In our implementation, index selection and the head GEMM are fused into a single custom CUDA kernel to avoid intermediate buffers and redundant global-memory traffic: the kernel performs warp-level column gathers with asynchronous, double-buffered shared-memory tiling, computes on Tensor Cores, and uses a persistent L2 access-policy window to keep activations cached. Verification is unchanged from standard speculative decoding.
5 Experimental Results
5.1 Experimental Settings
We adhere to the experimental setup and evaluation of the EAGLE (Li et al., 2024a, b) family and FR-Spec (Zhao et al., 2025). We implement the baseline FR-Spec and DynaSpec in the EAGLE 2 framework.
Datasets and Models. We benchmark DynaSpec over a diverse mix of seven tasks. Spec-Bench (Xia et al., 2024) supplies six of them—Machine Translation (WMT14 DE to EN) (Bojar et al., 2014), Multi-turn Conversation (MT-Bench) (Zheng et al., 2023), RAG and QA from Natural Questions (Kwiatkowski et al., 2019), Math from GSM8K (Cobbe et al., 2021), and Summarization (CNN/DailyMail) (Nallapati et al., 2016)—each evaluated on 80 prompts. The seventh task is Code, for which we employ HumanEval (Chen et al., 2021) with 164 problems. We bound generations to $1024$ tokens for Spec-Bench subtasks and HumanEval. We used Llama series of models, specifically Llama-33-8B-Instruct, which uses a vocabulary size of $128K$ for our experiments.
Evaluation metrics and Hardware Settings. We evaluate using metrics: #Mean Acceptance Length, the average number of tokens committed per draft–verification cycle. #Mean Acceptance Length is hardware-agnostic but does not capture the overhead of drafting. Our primary baseline is FR-Spec (Zhao et al., 2025), a static subset–selection approach built on EAGLE-2 (Li et al., 2024b); in contrast, DynaSpec performs dynamic subset selection atop EAGLE-2 (Li et al., 2024b). All experiments were run on a single NVIDIA A6000 GPU.
5.2 Mean Acceptance Length
Table 1 presents a comparison between DynaSpec and FR-Spec. We observe that DynaSpec achieves a higher mean acceptance length than FR-Spec while requiring a smaller mean shortlist size across seven datasets. Consistent with the findings of EAGLE-2, we also note that QA and summarization tasks yield lower accepted lengths compared to other tasks.
Config MT Conv. RAG Math QA Summ. Code Avg. Full Vocab $3.66$ $4.11$ $4.03$ $4.31$ $3.45$ $3.68$ $4.77$ $4.00$ vocab size=128k FR-Spec $3.38$ $3.87$ $3.85$ $4.16$ $3.32$ $3.51$ $4.11$ $3.74$ vocab size=32k DynaSpec $3.51$ $4.05$ $3.91$ $4.21$ $3.40$ $3.51$ $4.71$ $3.90$ vocab size 27,817 27,020 27,546 27,436 27,208 27,107 27,277 27,344
Table 1: Mean accepted tokens for Llama-3-8B on seven datasets. Vocab size indicates the shortlisted vocabulary size. Full Vocab (128k) corresponds to the EAGLE-2 baseline; FR-Spec uses 32k. For Dyna-Spec, shortlist size varies by dataset; per-dataset means are shown individually.
5.3 Position-Aware Vocabulary Shortlisting
Figure 2: Mean accepted length comparison for fixed top-$k$ clusters represented by DynaSpec-F vs. Position-aware top-$k$ clusters represented by DynaSpec-PA across seven datasets using Llama-3-8B model. Mean vocab size represents the mean shortlisted vocab size.
For the position-aware strategy, the results are presented in Figure 2. It shows that using a position-aware subset size outperforms a fixed size across all draft steps, as the first few tokens are more important than the later ones—an observation consistent with Gumiho Li et al. (2025). Table 2 further reports the results of combining the position-aware strategy with FR-Spec, which also demonstrates the effectiveness of this approach.
6 Conclusion
We address a central bottleneck in speculative decoding for large-vocabulary LLMs—the drafter’s $\mathcal{O}(|\mathcal{V}|d)$ output head. DynaSpec replaces static shortlists with a context-aware dynamic head that routes each prefix to a small number of token clusters and evaluates only over their union, while verification remains exact on the full vocabulary. Our analysis shows that context-conditioned supports strictly dominate any fixed subset in expected acceptance, and BCE-with-top-$k$ routing maximizes a lower bound on retained target mass. Systems-wise, we fuse index selection with the head GEMM and run routing in parallel, reducing draft-time cost. Empirically, DynaSpec increases mean acceptance length over fixed-shortlist baselines across tasks and models, without altering the target distribution. The method is plug-in compatible with EAGLE-style pipelines and orthogonal to cache reuse and serving-stack optimizations.
Appendix A Appendix
A.1 Position-aware frequency-ranked subset (PA-FR)
We ask whether position importance can also help semi-static, frequency-ranked shortlists. Let $\pi_{f}:\mathcal{V}\!\to\!\{1,\ldots,|\mathcal{V}|\}$ be the rank order induced by corpus token frequencies (lower is more frequent). Define a position-dependent budget $K_{\mathrm{fr}}(t)$ and construct a shortlist by taking the top-$K_{\mathrm{fr}}(t)$ most frequent tokens:
$K_{\mathrm{fr}}(t)=\begin{cases}k_{\max},&t\in\{0,1\},\\[4.0pt] \left\lfloor\dfrac{k_{\max}}{(t+1)}\right\rfloor,&t\geq 2.\end{cases}$
This position-aware schedule allocates a larger subset to early tokens, and decays the subset as generation proceeds, reducing head compute without retraining. Compared to a fixed frequency shortlist, PA-FR tends to improve mean accepted length at early steps while lowering average draft-time FLOPs, since the per-step head multiply scales with $K_{\mathrm{fr}}(t)$ rather than a constant $K$.
Config MT Conv. RAG Math QA Summ. Code Avg. FR-Spec-F $3.38$ $3.87$ $3.85$ $4.16$ $3.32$ $3.51$ $4.11$ $3.74$ vocab size = 32,768 FR-Spec-PA 3.50 3.94 3.86 4.18 3.38 3.58 4.21 3.81 vocab size = 31,739
Table 2: Mean accepted length comparison for Fixed top-$k$ tokens represented by FR-Spec-F vs. Position-aware top-$k$ tokens represented by FR-Spec-PA across seven datasets using Llama-3-8B model. Vocab size represents the mean shortlisted vocab size.
A.2 Meta classifier training
To train the meta classifier, we run eagle-2 Li et al. [2024b] on subsets of ShareGPT and UltraChat200K Ding et al. [2023], and collect the hidden states of the draft model, token embeddings, and the top-k tokens at each draft step. Using these, we construct the training dataset for the meta-classifier as $\{(h_{i-1},embed(x_{i})),map2meta(top\_k\_tokens_{i})\}$, where $h_{i}$ is the hidden state of the draft model at step $i$, $embed(x_{i})$ is the embedding of token $x_{i}$, and $map2meta(top\_k\_tokens_{i})$ represents the cluster IDs obtained by mapping the top-k tokens to their clusters.